OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 18 when x = 4. What is the constant of variation in decimal form?

OpenStudy (anonymous):

@igreen

OpenStudy (anonymous):

As \(y\) varies directly with \(x\), so you have : \[ y= kx\]

OpenStudy (anonymous):

Now, you have \(y=18\) and \(x=4\), so just plug the values in and find \(k\) there..

OpenStudy (anonymous):

ok so I do 4 . 18

OpenStudy (anonymous):

\[18 = k \cdot 4 \implies 18 = 4 \times k\] Divide by \(4\) both the sides to find \(k\) here..

OpenStudy (igreen):

You divide 4 to both sides.. 18 / 4 = ? @Im_a_BANANA

OpenStudy (anonymous):

4.5 @iGreen

OpenStudy (anonymous):

Yes, you are right.. :) Good..

OpenStudy (anonymous):

so 4.5 is the answer wow I feel stupid